I need help with 20-22Use the following to answer questions 20 – 22 Determine ending retained earnings for years 1, 2 and 3. At the beginning of ye
Answer #1

Ending Retained Earnings Balance = beginning Retained Earnings + Net income -Dividends

Year 1 = Ending Retained Earnings = 0+50,000-10,000 = 40,000 (Answer)

Year 2= Enging Retained Earnings = 40,000+85,000-50,000= 75,000(Answer)

Year 3= Ending Retained Earnings = 75,000+145,000-100,000 = 120,000(Answer)

Answers: Retained Earnings

20) $40,000

21) $75,000

22) $120,000